# CrossFi: A Cross Domain Wi-Fi Sensing Framework Based on Siamese Network

### v2

@@ VI Conclusion
This paper introduces CrossFi, a siamese network-based framework for cross-domain Wi-Fi sensing tasks utilizing CSI. In comparison to traditional siamese networks, we have designed a suitable framework to cater to different scenarios, including in-domain, few-shot, zero-shot cross-domain scenarios, and new-class scenario. Additionally, we have proposed an innovative attention-based method for similarity computation and an adaptively generated template method for the siamese network, which hold potential value for other machine learning domains as well. Through experiments conducted on Wi-Fi gesture recognition and people identification tasks, we have demonstrated that CrossFi achieves superior performance across different scenarios. As our approach is not specifically tailored for Wi-Fi sensing tasks, in the next step, we will evaluate its efficiency in other fields. Furthermore, we have observed that most models exhibit instability in cross-category scenarios. To ensure practical applicability, further improvements are necessary to enhance the model’s stability.

### v4

@@ VIII Conclusion
This paper introduces CrossFi, a siamese network-based framework for cross-domain Wi-Fi sensing tasks utilizing CSI. In comparison to traditional siamese networks, we have designed a suitable framework to cater to different scenarios, including in-domain, few-shot, zero-shot cross-domain scenarios, and new-class scenario. Additionally, we have proposed an innovative attention-based method for similarity computation and an adaptively generated template method for the siamese network, which hold potential value for other machine learning domains as well. Through experiments conducted on Wi-Fi gesture recognition and people identification tasks, we have demonstrated that CrossFi achieves superior performance across different scenarios. As our approach is not specifically tailored for Wi-Fi sensing tasks, in the next step, we will evaluate its efficiency in additional fields. Furthermore, we have observed that most models exhibit instability in cross-category scenarios. To ensure practical applicability, further improvements are necessary to enhance the model’s stability.
Biography
[Uncaptioned image] Zijian Zhao received the B.Eng. degree in computer science and technology from Sun Yat-sen University in 2024. He is currently pursuing the Ph.D. degree with Department of Civil and Environmental Engineering, The Hong Kong University of Science and Technology. He was a visiting student in Shenzhen Research Institute of Big Data from 2023 to 2024. His current research interests include intelligent transportation, wireless sensing, multi-agent reinforcement learning, and deep learning.
[Uncaptioned image] Tingwei Chen is currently a visiting student at the Shenzhen Research Institute of Big Data. He received his M.Sc. degree in communication engineering from The Chinese University of Hong Kong, Shenzhen, in 2023 and his B.S. degree in electronic information science and technology from Sun Yat-sen University in 2021. His research interests include multimodal machine learning and wireless sensing.
[Uncaptioned image] Zhijie Cai received the B.Sc. degree in mathematics and applied mathematics from Sun Yat-sen University in 2022. He is currently pursuing the Ph.D. degree with the Shenzhen Research Institute of Big Data, and the School of Science and Engineering, The Chinese University of Hong Kong, Shenzhen. His current research interests include edge intelligence, wireless sensing, and deep learning.
[Uncaptioned image] Xiaoyang Li (Member, IEEE) received the B.Eng. degree from the Southern University of Science and Technology (SUSTech) in 2016 and the Ph.D. degree from The University of Hong Kong in 2020. From 2020 to 2022, he was a Presidential Distinguished Research Fellow at SUSTech. He is currently an Associate Research Fellow with the Shenzhen Research Institute of Big Data, and an Adjunct Assistant Professor with the Chinese University of Hong Kong, Shenzhen. He is a recipient of the Youth Talent Support Program by China Communication Society, Forbes China 30 under 30, Overseas Youth Talent in Guangdong, Overseas High-caliber Personnel in Shenzhen, Outstanding Research Fellow in Shenzhen, the Best Paper Award of IEEE 4th International Symposium on Joint Communications & Sensing, and the Exemplary Reviewer of IEEE Wireless Communications Letters. He served as the Editor of Journal of Information and Intelligence, the Workshop/Session Chairs of IEEE ICASSP/ICC/WCNC/PIMRC/MIIS. His research interests include integrated sensing-communication-computation, edge learning, and network optimization.
[Uncaptioned image] Hang Li received the B.E. and M.S. degrees from Beihang University, Beijing, China, in 2008 and 2011, respectively, and the Ph.D. degree from Texas A&M University, College Station, TX, USA, in 2016. He was a postdoctoral research associate with Texas A&M and University of California-Davis (Sept. 2016-Mar. 2018). After being a visiting research scholar (Apr. 2018 – June 2019) at Shenzhen Research Institute of Big Data, Shenzhen, China, he has been a research scientist since June 2019. His current research interests include wireless networks, Internet of things, stochastic optimization, and applications of machine learning. He is recognized as Overseas High-Caliber Personnel (Level C) at Shenzhen in 2020.
[Uncaptioned image] Qimei Chen (Member, IEEE) received the Ph.D. degree from the College of Information Science and Electronic Engineering, Zhejiang University, Hangzhou, China, in 2017. She was a Visiting Student with the Department of Electrical and Computer Engineering, University of California at Davis, Davis, CA, USA, from 2015 to 2016. From 2017 to 2022, she was an Associate Researcher with the School of Electric Information, Wuhan University, Wuhan, China, where she has been an Associate Professor, since 2023. Her research interests include intelligent edge communication, unlicensed spectrum, massive MIMO-NOMA, and machine learning in wireless communications. She received the Exemplary Reviewer Certificate of the IEEE Wireless Communications Letters in 2020 and 2023. She has served as a workshop co-chair and TPC Member for IEEE conferences, such as ICC, GLOBECOM, PIMRC, and WCNC. She also served as a Guest Editor for the Special Issues on Heterogeneous Networks of Sensors and NOMA in ISAC (MDPI).
[Uncaptioned image] Guangxu Zhu (Member, IEEE) received the Ph.D. degree in electrical and electronic engineering from The University of Hong Kong in 2019. Currently he is a senior research scientist and deputy director of network system optimization center at the Shenzhen research institute of big data, and an adjunct associate professor with the Chinese University of Hong Kong, Shenzhen. His recent research interests include edge intelligence, semantic communications, and integrated sensing and communication. He is a recipient of the 2023 IEEE ComSoc Asia-Pacific Best Young Researcher Award and Outstanding Paper Award, the World’s Top 2% Scientists by Stanford University, the ”AI 2000 Most Influential Scholar Award Honorable Mention”, the Young Scientist Award from UCOM 2023, the Best Paper Award from WCSP 2013 and IEEE JSnC 2024. He serves as associate editors at top-tier journals in IEEE, including IEEE TMC, TWC and WCL. He is the vice co-chair of the IEEE ComSoc Asia-Pacific Board Young Professionals Committee.